OpenStudy (anonymous):

How do I solve this equation: 2^(4x) = 9^(x-1) for x?

OpenStudy (anonymous):

@satellite73 do you know how I can solve this?

OpenStudy (anonymous):

start with \[4x\ln(2)=(x-1)\ln(9)\] and then do some algebra

OpenStudy (anonymous):

I got up to: 4x = (x-1)log9/log2

OpenStudy (anonymous):

do I have to use natural log?

OpenStudy (anonymous):

Does not matter. Same answer.

OpenStudy (anonymous):

don't forget \(\ln(2)\) and \(\ln(9)\) are just numbers (constants) so treat them as such no you can use any log you like, but you only have two on your calculator

OpenStudy (anonymous):

but how would I solve this equation with log?

OpenStudy (anonymous):

think of them as \(a\) and \(b\)

OpenStudy (anonymous):

Okay... I guess I will let sattelite handle it :P .

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

\[4x\ln(2)=(x-1)\ln(9)\] \[4x\ln(2)=x\ln(9)-\ln(9)\] \[4\ln(2)x-\ln(9)x=-\ln(9)\] \[(4\ln(2)-\ln(9))x=-\ln(9)\] \[x=\frac{-\ln(9)}{4\ln(2)-\ln(9)}\] it is algebra all the way, and you should check my work because my algebra is lousy